Prove that |x+y+z|<=|x|+|y|+|z|


1
Expert's answer
2021-01-25T03:03:21-0500

"\text {let }a,b\in{R}"

"a*b \le|a|*|b|"

"(a+b)^2=a^2+b^2+2a*b\le{a^2+b^2+2|a|*|b|}"

"a^2+b^2+2|a|*|b|=|a|^2+|b|^2+2|a|*|b|=(|a|+|b|)^2"

"(a+b)^2\le(|a|+|b|)^2"

"\sqrt{(a+b)^2}\le\sqrt{(|a|+|b|)^2}"

"|a+b|\le|a|+|b|\ (1)"

"|x+y+z| \text{ let }a=x+y;b=z"

"\text{from inequality }(1)"

"|x+y+z|\le|x+y|+|z|\le|x|+|y|+|z|"

proved.